Solve 4p^2+4p=15 using the quadratic formula. Round to the nearest tenth if neccessary.
Answer by mananth(16946)

4p^2+4p=15
4p^2+4p-15=0
4p^2+4p+1 -16=0
(2p+1)^2=16
Take the square root on both sides
2p+1 = +/-4
2p= +4-1
2p=3
p=3/2
............
2p+1=-4
2p=-5
p=-5/2
p=3/2 OR -5/2
m.ananth@hotmail.ca